OpenStudy (anonymous):

10x + 8 + 5x - 4 = 12x + 16 Whats x value? PLEASE HELP.

OpenStudy (baby_bear69):

Can you combine like terms?

OpenStudy (anonymous):

I have no idea, im no good at math. I just need any value of x.

OpenStudy (baby_bear69):

10x+8+5x-4= 12x+16 15x+8-4 = 12x+16 15x+4 = 12x+16 -4 -4 15x = 12x+12 -12x -12x 3x = 12 -- -- 3 3 x=4

OpenStudy (baby_bear69):

Do you understand that?

OpenStudy (anonymous):

It says the answer is either 16, 40, 48, or 64. It's not 40, i already tried.

OpenStudy (anonymous):

Did you maybe copy the equation wrong onto here? @Baby_Bear69 's math is correct.

OpenStudy (baby_bear69):

That's a big difference o.o